# A Neural Material Point Method for Particle-based Simulations

### v2

@@ Abstract
Mesh-free Lagrangian methods are widely used for simulating fluids, solids, and their complex interactions due to their ability to handle large deformations and topological changes. These physics simulators, however, require substantial computational resources for accurate simulations. To address these issues, deep learning emulators promise faster and scalable simulations, yet they often remain expensive and difficult to train, limiting their practical use. Inspired by the Material Point Method (MPM), we present NeuralMPM, a neural emulation framework for particle-based simulations. NeuralMPM interpolates Lagrangian particles onto a fixed-size grid, computes updates on grid nodes using image-to-image neural networks, and interpolates back to the particles. Similarly to MPM, NeuralMPM benefits from the regular voxelized representation to simplify the computation of the state dynamics, while avoiding the drawbacks of mesh-based Eulerian methods. We demonstrate the advantages of NeuralMPM on several datasets, including fluid dynamics and fluid-solid interactions. Compared to existing methods, NeuralMPM reduces training times from days to hours, while achieving comparable or superior long-term accuracy, making it a promising approach for practical forward and inverse problems. A project page is available at https://neuralmpm.isach.be
@@ 1 Introduction
The Navier-Stokes equations describe the time evolution of fluids and their interactions with solid materials. As analytical solutions rarely exist, numerical methods are required to approximate the solutions. On the one hand, Eulerian methods discretize the fluid domain on a fixed grid, simplifying the computation of the dynamics, but requiring high-resolution meshes to solve small-scale details in the flow. Lagrangian methods, on the other hand, represent the fluid as virtual moving particles defining the system’s state, hence maintaining a high level of detail in regions of high density. While effective at handling deformations and topological changes (Monaghan, 2012), Lagrangian methods struggle with collisions and interactions with rigid objects (Vacondio et al., 2021, Lind et al., 2020).
@@
Regardless of the discretization strategy, large-scale high-resolution numerical simulations are computationally expensive, limiting their practical use in downstream tasks such as forecasting, inverse problems, or computational design. To address these issues, deep learning emulators have shown promise in accelerating simulations by learning an emulator model that can predict the system’s state at a fraction of the cost. Next to their speed, neural emulators also have the strategic advantage of being differentiable, enabling their use in inverse problems and optimization tasks (Allen et al., 2022, Zhao et al., 2022, Forte et al., 2022). Moreover, they have the potential to be learned directly from real data, bypassing the costly and resource-intensive process of building a simulator (Lam et al., 2023, Pfaff et al., 2021, Lemos et al., 2023, Jumper et al., 2021, He et al., 2019). In this direction, particle-based neural emulators (Sanchez-Gonzalez et al., 2020, Ummenhofer et al., 2020, Prantl et al., 2022) have seen success in accurately simulating fluids and generalizing to unseen environments. These emulators, however, suffer from the same issues as traditional Lagrangian methods, with collisions and interactions with rigid objects being particularly challenging. These emulators may also require long training and inference times, limiting their practical use.
Taking inspiration from the hybrid Material Point Method (MPM) (Sulsky et al., 1993, Nguyen et al., 2023) that combines the strengths of both Eulerian and Lagrangian methods, we introduce NeuralMPM, a neural emulation framework for particle-based simulations. As in MPM, NeuralMPM maintains Lagrangian particles to represent the system’s state but models the system dynamics on voxelized representations. In this way, NeuralMPM benefits from a regular grid structure to simplify the computation of the state dynamics but avoids the drawbacks of mesh-based Eulerian methods. By interpolating the particles onto a fixed-size grid, it also bypasses the need to perform an expensive neighbor search at every timestep, substituting it with two interpolation steps based on cheap voxelization (Xu et al., 2021). By defining the system dynamics on a grid, NeuralMPM can also leverage well-established grid-to-grid neural architectures. The resulting inductive bias allows the model to more easily process the global and local structures of the point cloud, instead of having to discover them, and frees capacity for learning the dynamics of the system represented by the grid. Compared to previous data-driven approaches (Sanchez-Gonzalez et al., 2020, Ummenhofer et al., 2020, Prantl et al., 2022), these improvements reduce the training time from days to hours, while achieving higher or comparable accuracy.
@@ 2 Computational Fluid Dynamics
Computational fluid dynamics simulations can be classified into two broad categories, Eulerian and Lagrangian, depending on the discretization of the fluid (Rakhsha et al., 2021). In Eulerian simulations, the domain is discretized with a mesh, with state variables $u^{t}_{i}$ (such as mass or momentum) maintained at each mesh point $i$. Well-known examples of Eulerian simulations are the finite difference method, where the domain is divided into a uniform regular grid (also called an Eulerian grid), and the finite element method, where the domain is divided into regions, or elements, that may have different shapes and density, allowing to increase the resolution in only some areas of the domain (Iserles, 2008, Morton & Mayers, 2005). Lagrangian simulations, on the other hand, discretize the fluid as a set of virtual moving particles $\{p^{t}_{i},u^{t}_{i}\}_{i=1}^{N}$, each described by its position $p^{t}_{i}$ and state variables $u^{t}_{i}$ that include the particle velocity $v^{t}_{i}$. To simulate the fluid, the particles move according to the dynamics of the system, producing a new set of particles $\{p^{t+1}_{i},u^{t+1}_{i}\}_{i=1}^{N}$ at each timestep. Simulations in Lagrangian coordinates are particularly useful when the fluid is highly deformable, as the particles can move freely and adapt to the fluid’s shape. Among Lagrangian methods, Smoothed Particle Hydrodynamics (SPH) is one of the most popular, where the fluid is represented by a set of particles that interact with each other through a kernel function that smooths the interactions. SPH has been widely used in large-scale astrophysical simulations, such as galactic dynamics (Wissing & Shen, 2023, Few et al., 2016) or planetary collision (Kegerreis et al., 2022), and in ocean engineering (Gotoh & Khayyer, 2016, Tan et al., 2023, Lyu et al., 2022) to model deformations and fractures.
@@
Notwithstanding the success of numerical simulators, they remain expensive, slow, and, most of the time, non-differentiable. In recent years, differentiable neural emulators have shown great promise in accelerating fluid simulations, most notably in a series of works to emulate SPH simulations in a fully data-driven manner. Graph network-based simulators (GNS) (Sanchez-Gonzalez et al., 2020) use a graph neural network (GNN) and a graph built from the local neighborhood of the particles to predict the acceleration of the system. The approach requires building a graph out of the point cloud at every timestep to obtain structural information about the cloud, which is an expensive operation. In addition, the GNN needs to extract global information from its nodes, which is only possible with a high number of message-passing steps, resulting in a large computational graph and long training and inference times. This large computational graph, along with repeated construction, makes fully autoregressive training over long rollouts impractical, as the gradients need to backpropagate all the way back to the first step. Cheaper strategies exist, like the push-forward trick (Brandstetter et al., 2022), but they have been shown to be inferior to fully backpropagating through trajectories (List et al., 2024, Sharabi & Louppe, 2023). As autoregressive training is not available, the stability of the learned dynamics can be compromised, making the model prone to diverging or oscillating. Noise injection training strategies can be used to increase the stability of the rollouts, but the magnitude of the noise becomes a critical parameter. Han et al. (2022) introduce improvements to GNS to make them subequivariant to certain transformations. They show increased accuracy on simulations involving solid objects. An alternative approach is the continuous convolution (CConv) (Ummenhofer et al., 2020, Winchenbach & Thuerey, 2024), an extension of convolutional networks to point clouds. In this method, a convolutional kernel is applied to each particle by interpolating the values of the kernel at the positions of its neighbors, which are found via spatial hashing on GPU, a cheaper alternative to tree-based searches that allows for autoregressive training. In (Prantl et al., 2022), Deep Momentum Conserving Fluids (DMCF) build upon CConv to design a momentum-conserving architecture. Nevertheless, to account for long-range interactions, the authors introduce different branches, with different receptive fields, into their network. The number of branches, and their hyperparameters, need to be tuned to capture global dependencies, leading to long training times even with optimized CUDA kernels. Finally, Zhang et al. (2020), propose an approach that uses nearest neighbors to construct the local features of each particle. Those local features are then averaged onto a regular grid. Like GNS, this method suffers from the need to repeat the neighbor search at every simulation timestep. Ultimately, the performance of point cloud-based simulators is tightly linked to the method used to process the spatial structure of the cloud. Brute force neighbour search is $\mathcal{O}(N^{2})$, K-d trees are $\mathcal{O}(N\log{N})$, and voxelization and hashing are $\mathcal{O}(N)$ (Xu et al., 2021, Hastings & Mesit, 2005).
@@
An alternative to data-driven modeling is the use of hybrid models, where parts of a classical solver are replaced with learned components. For instance, Yin et al. (2021) employ a neural network to learn unknown physics, which is then integrated into a simulator. Similarly, Li et al. (2024) use a neural network to bypass computational bottlenecks in MPM simulators, while Ma et al. (2023) learn general constitutive laws, allowing for one-shot trajectory learning. These approaches achieve impressive results by leveraging extensive physics knowledge, but this reliance also limits their applicability. Hybrid models inherit both the strengths and weaknesses of classical and ML methods.
@@ 3 NeuralMPM
We consider a Lagrangian system evolving in time and defined by the positions $p^{t}_{i}$ and velocities $v^{t}_{i}$ of a set of $N$ particles $i=1,...,N$. We denote with $P^{t}$ and $V^{t}$ the set of positions and velocities of all particles at time $t$ and with $S^{t}=(P^{t},V^{t})$ the full state of the system. In a more complex setting, the state of the system can include other local properties, such as pressure or elastic stiffness of materials, and global properties, such as an external force. In this work, for simplicity, we let the network learn the relevant simulation parameters implicitly. The evolution of the particles is described by a function $f$ mapping the current state of the system to its next state $S^{t+1}=f(S^{t})$. Given a starting system $S^{0}=(P^{0},V^{0})$, its full trajectory, or rollout, is denoted by $S^{1:T}$. Our goal is to build an emulator $\hat{f}_{\theta}(\cdot)$ capable of predicting a full rollout $\hat{f}_{\theta}^{1:T}(S^{0})$ of $T$ timesteps from the initial state $S^{0}$. Following MPM, NeuralMPM operates in four steps, as illustrated in Figure 1:
@@
Taking advantage of the regular grid representation of the cloud, the grid velocities $\{\hat{V}^{i}\}_{i=t+1}^{t+m}$ of the next $m$ timesteps are predicted using a neural network. We chose a U-Net (Ronneberger et al., 2015) as it is a well-established image-to-image model, known to perform well in various tasks, including physical applications. The combination of kernels applied with different receptive fields (from smaller to larger) allows the U-Net to efficiently extract both local and global information. Nonetheless, any grid-to-grid architecture could be used. We experiment with the FNO (Li et al., 2021) architecture in Appendix B and find it to underperform, leading us to keep the U-Net. A fully convolutional U-Net and an FNO have the additional advantage of being able to generalize to different domain shapes, a desirable property (Section 4.3).
@@
NeuralMPM is trained end-to-end on a set of trajectories $S^{0:T}$ to minimize the mean squared error $||P^{t+1}-\hat{P}_{\theta}^{t+1}(S^{t})||^{2}_{2}$ between the ground-truth and predicted next positions of the particles. At inference time, the model is exposed to much longer sequences, which requires carefully stabilizing the rollout procedure to prevent the accumulation of large errors over time. To address this, we first make use of autoregressive training (Prantl et al., 2022, Ummenhofer et al., 2020), where the model is unrolled $K$ times on its own predictions, producing a sequence of $\hat{S}^{k}=\hat{f}_{\theta}(\hat{S}^{k-1})$ for $k=1,...,K$ and initial input $\hat{S}^{0}=S^{0}$, before backpropagating the error through the entire rollout. Unlike more costly methods that require alternative stabilization strategies, such as noise injection (Sanchez-Gonzalez et al., 2020), NeuralMPM’s efficiency makes autoregressive training possible. Nevertheless, to further stabilize the training, we couple autoregressive training with time bundling (Brandstetter et al., 2022), resulting in a training strategy where the model predicts $m$ steps $\hat{S}^{1:m}$ at once from a single initial state, inside an outer autoregressive loop of $K$ steps of length $m$. We show in Section 4 that this training strategy leads to more accurate rollouts.

### v3

# A Neural Material Point Method for Particle-based Emulation

@@ Abstract
Mesh-free Lagrangian methods are widely used for simulating fluids, solids, and their complex interactions due to their ability to handle large deformations and topological changes. These physics simulators, however, require substantial computational resources for accurate simulations. To address these issues, deep learning emulators promise faster and scalable simulations, yet they often remain expensive and difficult to train, limiting their practical use. Inspired by the Material Point Method (MPM), we present NeuralMPM, a neural emulation framework for particle-based simulations. NeuralMPM interpolates Lagrangian particles onto a fixed-size grid, computes updates on grid nodes using image-to-image neural networks, and interpolates back to the particles. Similarly to MPM, NeuralMPM benefits from the regular voxelized representation to simplify the computation of the state dynamics, while avoiding the drawbacks of mesh-based Eulerian methods. We demonstrate the advantages of NeuralMPM on 6 datasets, including fluid dynamics and fluid-solid interactions simulated with MPM and Smoothed Particles Hydrodynamics (SPH). Compared to GNS and DMCF, NeuralMPM reduces training time from 10 days to 15 hours, memory consumption by 10x-100x, and increases inference speed by 5x-10x, while achieving comparable or superior long-term accuracy, making it a promising approach for practical forward and inverse problems. A project page is available at https://neuralmpm.isach.be.
@@ 1 Introduction
The Navier-Stokes equations describe the time evolution of fluids and their interactions with solid materials. As analytical solutions rarely exist, numerical methods are required to approximate the solutions. On the one hand, Eulerian methods discretize the fluid domain on a fixed grid, simplifying the computation of the dynamics, but requiring high-resolution meshes to solve small-scale details in the flow. Lagrangian methods, on the other hand, represent the fluid as virtual moving particles defining the system’s state, hence maintaining a high level of detail in regions of high density. While effective at handling deformations and topological changes (Monaghan, 2012), Lagrangian methods struggle with collisions and interactions with rigid objects (Vacondio et al., 2021; Lind et al., 2020).
@@
Regardless of the discretization strategy, large-scale high-resolution numerical simulations are computationally expensive, limiting their practical use in downstream tasks such as forecasting, inverse problems, or computational design. To address these issues, deep learning emulators have shown promise in accelerating simulations by learning an emulator model that can predict the system’s state at a fraction of the cost. Next to their speed, neural emulators also have the strategic advantage of being differentiable, enabling their use in inverse problems and optimization tasks (Allen et al., 2022; Zhao et al., 2022; Forte et al., 2022). Moreover, they have the potential to be learned directly from real data, bypassing the costly and resource-intensive process of building a simulator (Lam et al., 2023; Pfaff et al., 2021; Lemos et al., 2023; Jumper et al., 2021; He et al., 2019). In this direction, particle-based neural emulators (Sanchez-Gonzalez et al., 2020; Ummenhofer et al., 2020; Prantl et al., 2022) have seen success in accurately simulating fluids and generalizing to unseen environments. These emulators, however, suffer from the same issues as traditional Lagrangian methods, with collisions and interactions with rigid objects being particularly challenging. These emulators can also require long training and inference times, limiting their practical use.
Taking inspiration from the hybrid Material Point Method (MPM) (Sulsky et al., 1993; Nguyen et al., 2023) that combines the strengths of both Eulerian and Lagrangian methods, we introduce NeuralMPM, a neural emulation framework for particle-based simulations.
  • As in MPM, NeuralMPM uses Lagrangian particles to represent the system’s state but models system dynamics on voxelized grids. This approach benefits from a regular grid structure for simplified state dynamics computation while avoiding the drawbacks of mesh-based Eulerian methods.
  • It interpolates particles onto a fixed-size grid, bypassing the need for expensive neighbor searches at each timestep and replacing them with two efficient voxelization-based interpolation steps (Xu et al., 2021)
  • By defining dynamics on a grid, NeuralMPM leverages well-established grid-to-grid neural architectures. This introduces an inductive bias that enables easier processing of global and local point cloud structures, freeing capacity to learn the system’s dynamics.
Compared to previous data-driven approaches (Sanchez-Gonzalez et al., 2020; Ummenhofer et al., 2020; Prantl et al., 2022), these improvements reduce the training time from days to hours, while achieving higher or comparable accuracy.
@@ 2 Computational Fluid Dynamics
Computational fluid dynamics simulations can be classified into two broad categories, Eulerian and Lagrangian, depending on the discretization of the fluid (Rakhsha et al., 2021). In Eulerian simulations, the domain is discretized with a mesh, with state variables $u^{t}_{i}$ (such as mass or momentum) maintained at each mesh point $i$. Well-known examples of Eulerian simulations are the finite difference method, where the domain is divided into a uniform regular grid (also called an Eulerian grid), and the finite element method, where the domain is divided into regions, or elements, that may have different shapes and density, allowing to increase the resolution in only some areas of the domain (Iserles, 2008; Morton & Mayers, 2005). Another widely used family of Eulerian methods is the spectral and pseudo-spectral family of methods, where the equation is solved in Fourier space or alternating between Fourier and real space, respectively, leveraging the fact that differentiation in real space is multiplication in Fourier space Pope (2000). Lagrangian simulations, on the other hand, discretize the fluid as a set of virtual moving particles $\{p^{t}_{i},u^{t}_{i}\}_{i=1}^{N}$, each described by its position $p^{t}_{i}$ and state variables $u^{t}_{i}$ that include the particle velocity $v^{t}_{i}$. To simulate the fluid, the particles move according to the dynamics of the system, producing a new set of particles $\{p^{t+1}_{i},u^{t+1}_{i}\}_{i=1}^{N}$ at each timestep. Simulations in Lagrangian coordinates are particularly useful when the fluid is highly deformable, as the particles can move freely and adapt to the fluid’s shape. Among Lagrangian methods, Smoothed Particle Hydrodynamics (SPH) (Gingold & Monaghan, 1977; Price, 2012) is one of the most popular, where the fluid is represented by a set of particles that interact with each other through a kernel function that smooths the interactions.
@@
Notwithstanding the success of numerical simulators, they remain expensive, slow, and, most of the time, non-differentiable. In recent years, differentiable neural emulators have shown great promise in accelerating fluid simulations, most notably in a series of works to emulate SPH simulations in a fully data-driven manner. Graph network-based simulators (GNS) (Sanchez-Gonzalez et al., 2020) use a graph neural network (GNN) and a graph built from the local neighborhood of the particles to predict the acceleration of the system. The approach requires building a graph out of the point cloud at every timestep to obtain structural information about the cloud, which is an expensive operation. In addition, the GNN needs to extract global information from its nodes, which is only possible with a high number of message-passing steps, resulting in a large computational graph and long training and inference times. This large computational graph, along with repeated construction, makes fully autoregressive training over long rollouts impractical, as the gradients need to backpropagate all the way back to the first step. Cheaper strategies exist, like the push-forward trick (Brandstetter et al., 2022b), but they have been shown to be inferior to fully backpropagating through trajectories (List et al., 2024; Sharabi & Louppe, 2023). As autoregressive training is not available, the stability of the learned dynamics can be compromised, making the model prone to diverging or oscillating. Noise injection training strategies can be used to increase the stability of the rollouts, but the magnitude of the noise becomes a critical parameter. Related approaches to GNS include Han et al. (2022), who introduce improvements to GNS to make them subequivariant to certain transformations and show increased accuracy on simulations involving solid objects, and Viswanath et al. (2024), who apply GNS on a reduced representation of the particle system obtained by farthest point sampling.
An alternative approach is the continuous convolution (CConv) (Ummenhofer et al., 2020; Winchenbach & Thuerey, 2024), an extension of convolutional networks to point clouds. In this method, a convolutional kernel is applied to each particle by interpolating the values of the kernel at the positions of its neighbors, which are found via spatial hashing on GPU, a cheaper alternative to tree-based searches that allows for autoregressive training. In (Prantl et al., 2022), Deep Momentum Conserving Fluids (DMCF) build upon CConv to design a momentum-conserving architecture. Nevertheless, to account for long-range interactions, the authors introduce different branches, with different receptive fields, into their network. The number of branches, and their hyperparameters, need to be tuned to capture global dependencies, leading to long training times even with optimized CUDA kernels. Finally, Zhang et al. (2020) propose an approach that uses nearest neighbors to construct the local features of each particle. Those local features are then averaged onto a regular grid. Like GNS, this method suffers from the need to repeat the neighbor search at every simulation timestep. Ultimately, the performance of point cloud-based simulators is tightly linked to the method used to process the spatial structure of the cloud. Brute force neighbour search is $\mathcal{O}(N^{2})$, K-d trees are $\mathcal{O}(N\log{N})$, and voxelization and hashing are $\mathcal{O}(N)$ (Xu et al., 2021; Hastings & Mesit, 2005). Recently, Alkin et al. (2024), attempt to bridge the gap between Lagrangian and Eulerian representations by avoiding explicit grid- or particle-based latent structures, but necessitating extra networks to account for particle movement or empty volumes.
@@
An approach different from data-driven modeling is hybrid models, where parts of a classical solver are replaced with learned components. For instance, Yin et al. (2021) employ a neural network to learn unknown physics, which is then integrated into a simulator. Similarly, Li et al. (2024a) use a neural network to bypass computational bottlenecks in MPM simulators, while Ma et al. (2023) learn general constitutive laws, allowing for one-shot trajectory learning. These approaches achieve impressive results by leveraging extensive physics knowledge, but this reliance also limits their applicability. Hybrid models may inherit both the strengths and weaknesses of classical and ML methods. Our approach falls within the data-driven domain, as the prior constraints on the learnable dynamics induced by the MPM-like algorithm are minimal and can be overcome with data (Table 1).
@@ 3 NeuralMPM
We consider a Lagrangian system evolving in time and defined by the positions $p^{t}_{i}$ and velocities $v^{t}_{i}$ of a set of $N$ particles $i=1,...,N$. We denote with $P^{t}$ and $V^{t}$ the set of positions and velocities of all particles at time $t$ and with $S^{t}=(P^{t},V^{t})$ the full state of the system. In a more complex setting, the state of the system can include other local properties, such as pressure or elastic stiffness of materials, and global properties, such as an external force. In this work, for simplicity, we let the network learn the relevant simulation parameters implicitly. The evolution of the particles is described by a function $f$ mapping the current state of the system to its next state $S^{t+1}=f(S^{t})$. Given a starting system $S^{0}=(P^{0},V^{0})$, its full trajectory, or rollout, is denoted by $S^{1:T}$. Our goal is to build an emulator $\hat{f}_{\theta}(\cdot)$ capable of predicting a full rollout $\hat{f}_{\theta}^{1:T}(S^{0})$ of $T$ timesteps from the initial state $S^{0}$. Following MPM, NeuralMPM operates in four steps, described below, as illustrated in Figure 1 and in Alg 1.
@@
Taking advantage of the regular grid representation of the cloud, the grid velocities $\{\hat{V}_{g}^{i}\}_{i=t+1}^{t+m}$ of the next $m$ timesteps are predicted using a neural network. We chose a U-Net (Ronneberger et al., 2015) as it is a well-established image-to-image model, known to perform well in various tasks, including physical applications. The combination of kernels applied with different receptive fields (from smaller to larger) can allow the U-Net to efficiently extract both local and global information. Nonetheless, any grid-to-grid architecture could be used. We experiment with the FNO (Li et al., 2021) architecture in Appendix C and find it to underperform, leading us to keep the U-Net. A fully convolutional U-Net and an FNO have the additional advantage of being able to generalize to different domain shapes, a desirable property (Section 4.3).
@@
NeuralMPM is trained end-to-end on a set of trajectories $S^{0:T}$ to minimize the mean squared error $||P^{t+1}-\hat{P}_{\theta}^{t+1}(S^{t})||^{2}_{2}$ between the ground-truth and predicted next positions of the particles. At inference time, the model is exposed to much longer sequences, which requires carefully stabilizing the rollout procedure to prevent the accumulation of large errors over time. To address this, we first make use of autoregressive training (Prantl et al., 2022; Ummenhofer et al., 2020), where the model is unrolled $K$ times on its own predictions, producing a sequence of $\hat{S}^{k}=\hat{f}_{\theta}(\hat{S}^{k-1})$ for $k=1,...,K$ and initial input $\hat{S}^{0}=S^{0}$, before backpropagating the error through the entire rollout. Unlike more costly methods that require alternative stabilization strategies, such as noise injection (Sanchez-Gonzalez et al., 2020), NeuralMPM’s efficiency makes autoregressive training possible. Nevertheless, to further stabilize the training, we couple autoregressive training with time bundling (Brandstetter et al., 2022b), resulting in a training strategy where the model predicts $m$ steps $\hat{S}^{1:m}$ at once from a single initial state, inside an outer autoregressive loop of $K$ steps of length $m$. We show in Section 4 that this training strategy leads to more accurate rollouts.
Algorithm 1 NeuralMPM
1:$P^{0}$, $V^{0}$, grid $g$, $t$, neural backbone $h$, functions p2g and g2p
2:$\hat{P}^{t}\leftarrow P^{t}$, $\hat{V}^{t}\leftarrow V^{t}$
3:while $t<T$ do
4:     Voxelization: Interpolate density and velocities onto the grid using $G^{t}=(D_{g}^{t},V_{g}^{t})=\textbf{p2g}(\hat{P}^{t},\hat{V}^{t})$
5:     Processing: $\{\hat{V}_{g}^{i}\}_{i=t+1}^{t+m}=h(G^{t})$
6:     for $i$ in range($1,m$do
7:         Update velocities: $\hat{V}^{t+i}=\textbf{g2p}(\hat{V}_{g}^{t+i},\hat{P}^{t+i-1})$
8:         Update positions: $\hat{P}^{t+i}=P^{t+i-1}+\Delta t\hat{V}^{t+i}$
9:     end for
10:     $t\leftarrow t+m$
11:end while
12:return Full trajectory $\{\hat{P},\hat{V}\}_{t}^{T}$